# In the Network of the Conclave: Social Network Analysis and the Making of a Pope

### v1

@@ Abstract
This study brings a network perspective to the centuries-old mystery of papal elections by mapping the relational architecture of the College of Cardinals. Using publicly available data sources, such as official Vatican directories and episcopal consecration records, we assemble a multilayered network that captures cardinals’ co-membership in various collegial bodies of the Vatican and their consecration ties. We then calculate structural metrics to capture three key mechanisms that we suggest have a crucial role in the dynamics of the conclave: status, mediation power, and coalition building. Our descriptive study—publicly shared before the white smoke of May 8, 2025—uncovers a striking case: Cardinal Robert F. Prevost, largely ignored by pundits, bookmakers, and AI models, held a uniquely advantageous position in the Vatican network, by virtue of being both highly central and positioned in-between others. Thus, although being considered an underdog by many, the network perspective suggests that Cardinal Prevost (elected as Pope Leo XIV) was de facto one of the strongest “papabili.” In highlighting the divergence between network-agnostic and network-informed analyses, we show the merit of adopting a social network perspective to study the conclave. Rather than predicting ballots, our work demonstrates how network theory can reveal the structural preconditions for consensus. We argue that any serious inquiry into future conclaves—or analogous elite decision processes—must begin by charting the relational space of influence before considering doctrinal, regional, or theological dimensions.
Keywords: social network analysis centrality measures papal conclave
@@ 1 Introduction
A: “I would not say so, in the sense that the Holy Spirit picks out the Pope… I would say that the Spirit does not exactly take control of the affair, but rather, like a good educator, as it were, leaves us much space, much freedom, without entirely abandoning us. Thus the Spirit’s role should be understood in a much more elastic sense—not that he dictates the candidate for whom one must vote. Probably the only assurance he offers is that the thing cannot be totally ruined… There are too many contrary instances of popes the Holy Spirit obviously would not have picked!
(Interview with Cardinal Ratzinger, later Pope Benedictus XVI; emphasis ours)
Few decision-making processes in the modern world remain as cloaked in rituals, secrecy, and symbolism as the papal conclave. When the College of Cardinals gathers in the Sistine Chapel to elect a new pope, the world watches for the white smoke that signals the outcome of a process often described as guided by divine inspiration. Yet, as the quote above suggests, the process is not so much steered by divine forces as it is shaped by human judgment and negotiation. The very same existence of “contrary instances”—as Cardinal Ratzinger stated—reveals that the conclave, as many others human organizations, is not immune to mechanisms such as mediation power, coalition building, and status dynamics, which can play a crucial role in influencing the voting dynamics that determine the new pope of the Roman Catholic Church.
@@
Although the cardinals in the conclave are secluded from the rest of the world, they are not isolated from the force field of social influences among them built through interpersonal connections (Padgett and Ansell,, 1993; Krackhardt,, 1990). Indeed, we argue that the conclave is a complex, highly relational decision-making process, where influence flows through network ties, which form the social scaffolding through which consensus can be built (Urena et al.,, 2019; Stevenson and Greenberg,, 2000). These ties create a complex network that connects the cardinals, allowing them to get to know each other, work together, exchange ideas, and transfer information (Granovetter,, 1973; Reagans and McEvily,, 2003). Given this relational structure, therefore, we ask: What role does the network—and the structural position each cardinal occupies within it—play in shaping the likelihood that someone can emerge as frontrunner candidate during the conclave?
@@
The question is as consequential as difficult to answer. Obtaining reliable data to understand why some cardinals emerge as key figures and how the college converges toward a candidate—and, more generally, what happens during the conclave—is, in fact, extremely challenging, if not impossible. Those who participate in the conclave are bound to absolute silence, even years after the fact, under penalty of excommunication. [^1] The Apostolic Constitution Universi Dominici Gregis (John Paul II, 22 February 1996), which sets forth the norms for the conduct of the conclave, establishes the penalty of latae sententiae excommunication: “for those involved in the conclave who are bound to secrecy, in the event of violating that secrecy (n. 58)”. But does this mean we should abandon a scientific approach and resort to speculation? We argue not.
@@
In this paper, we advocate for applying social network theories and methods to enhance the understanding of the papal conclave by examining how interpersonal connections among cardinals may shape decision-making (Borgatti et al.,, 2009). This relational perspective allows us to move beyond anecdotal interpretations, offering a systematic lens for studying how patterns of connections may influence voting behavior. While certain ties—such as trust or support—may emerge or evolve during the conclave, the backbone of the network structure, built through long-standing institutional, personal, and professional relationships, originates prior to it. Analyzing the preexisting networks, therefore, provides a valuable vantage point for anticipating the dynamics of consensus formation within the conclave.
Specifically, while acknowledging that formal ecclesiastical hierarchies define roles and responsibilities, we suggest that influence and consensus dynamics within the conclave are embedded in relations, which can be made visible and analytically tractable through a network perspective (Pfeffer,, 2007). In doing so, we can model the conclave as a complex multiplex social system, and thus illuminate key network mechanisms underlying papal succession.
@@
Our approach is rooted entirely in network theory and mathematics, taking advantage of a decades of research on network status, information control and mediation power, and coalition building (Podolny,, 2001; Burt,, 1992; Mizruchi,, 1992). We did not leverage natural language processing, deep learning, or other artificial intelligence tools. Rather, this is a case of applying a classic social network analysis approach to a novel setting—leveraging co-membership ties and apostolic lineages—to shed light on one of the most secretive decision-making processes in the world. Our findings suggest that understanding who is papabile is not only a matter of ideology, but also of relational embeddedness, that is, being known, trusted, and connected in the right ways to the right others (Useem,, 1986; Heinz,, 1993).
@@ 2.1 Collegial Structures
Thus, while the hierarchical structure of the Catholic Church reflects a vertical chain of command, with the pope at the apex, such a depiction risks obscuring the complex, consultative, and collaboration-based processes that characterize its governance and organization. Central to this collaborative approach is the role of various collegial bodies, such as dicasteries, congregations, commissions, and councils, which support and advise the pope in exercising his office. These include institutions like the Dicastery for the Doctrine of the Faith, the Council for the Economy, and the Synod of Bishops, all of which function with a spirit of shared responsibility and collective discernment. [^4] Since the Second Vatican Council (1962–1965), there has been a significant theological and organizational emphasis on episcopal collegiality—the principle that, while the Pope holds primacy, bishops collectively share responsibility for governing the universal Church. Thus, while final authority rests with the pope, the actual process involves substantial interaction, negotiation, and consensus-building among several actors. Like in many large and complex organizations, also in the Catholic Church the extensive use of team- or group-based decision-making and coordination mechanisms, such as committees, task forces, or project teams, creates an articulated web of interactions among individuals by virtue of shared presence within these groups.
From a theoretical point of view, this approach has similarities with the idea of foci proposed by Feld, (1981: 1018) in that “Foci (…) may actively bring people together or passively constrain them to interact.” While Feld’s (1981) concept of foci is broader than the organizational focus studied here, we suggest that the formal dimensions of an organization are very powerful structural foci for intra-organizational interactions (Coser,, 1974; Merton,, 1976). This idea echoes Shipilov et al., (2014: 73) work on bureaucrats, proposing the idea of structured foci defined as “[…] formal entities that actively and regularly bring individuals together to engage in organized joint activities created specifically to encourage members to form personal bonds.” Similarly, Dahlander and McFarland, (2013) embrace the idea of formal organization as meaningful foci for its members. They demonstrate that individuals who share organizational foci are more likely to form social ties. Moreover, it is not simply the formation of ties that is shaped by organizational foci, but rather the evolutionary pattern of informal networks. For example, in the context of project organizations, Anjos and Reagans, argue that actors’ interproject mobility has a negative effect on network connections: “when two individuals stop working together, their network connection decays quickly” (Anjos and Reagans, 2020: 11).
@@
Similar to directors sitting in corporate boards (Mizruchi,, 1996), or actors collaborating in movies (Watts and Strogatz,, 1998), the shared participation of cardinals in the formal collegial bodies of the Roman Curia constitute ties in the co-membership networks (Breiger,, 1974). Thus, to capture and analyze how structural influence unfolds among cardinals, we used a well-established technique from social network analysis by projecting a bipartite network—linking cardinals to collegial organizational bodies—into a cardinal-by-cardinal co-membership network. While our context differs, we draw methodological inspiration from Porter et al., (2005), who applied a similar approach to study co-membership patterns among legislators in the U.S. House of Representatives. In our case, the two-mode network connects cardinals to the collegial bodies to which they are assigned. Through collaboration within these bodies, cardinals cultivate mutual familiarity, exchange ideas, information, and expertise, and provide reciprocal support. To summarize, the network-based lens highlights that decision-making in the Catholic Church—especially at the Vatican level—is shaped through complex webs of interlocking memberships and collaborative structures. As such, this network originating from formal assignments constitutes a fundamental component of the salient relational infrastructure relevant to the conclave.
@@
To achieve a more fine-grained understanding of the network that connects those entrusted with electing the pope, formal ties should be complemented by more informal connections—such as friendship, advice, or trust (Krackhardt et al.,, 1992). Although our study could not reliably capture informal relationships of this kind for all cardinals, we rigorously reconstructed another type of tie that enriches the co-membership structure and helps illuminate the multiple social spaces that characterize the conclave’s decision-making process. [^5] For a small subset of cardinals, we were able to supplement our data with journalistic reports suggesting the presence of informal ties (e.g., personal friendships or long-standing collaborations). We ran our analyses both with and without these additional ties and found no meaningful differences in the results. However, because these data are not systematically available across the full sample, we do not include them in the reported network. Specifically, we focus on lines of episcopal consecration, which we describe in detail in the following section.
@@ 2.2 Consecration Lines
In the context of the conclave, these ties take on a concrete political and relational significance. A cardinal who has consecrated another member of the college—whether as principal or co-consecrator—has played a formative role in that person’s ecclesiastical trajectory. This relationship may imply a sense of loyalty, spiritual indebtedness, or alignment of vision. As such, consecration lines create a relational infrastructure of influence and obligation within the electoral body, shaping how support, trust, and alignment might unfold during the papal election.
@@ 3 Network Mechanisms and the Conclave
Despite its highly codified procedures, the conclave is far from a static or predictable event. One key reason is that the actual voting preferences and dynamics remain entirely confidential, even after the end of the conclave. [^6] Allegedly, the only conclave on which there is information on the voting patterns over different rounds is the 2013 conclave that ended with the election of Cardinal Bergoglio as Pope Francis. This information comes from a book wrote by journalist Gerard O’Connel (O’Connell,, 2019). Ballots are burned, and no official record of individual votes is ever released. This secrecy makes each conclave a unique and analytically challenging decision-making episode. It reasonable to expect that electors may enter the conclave with a short list of preferred candidates—often shaped by theological affinity, regional interests, curial loyalties, pastoral vision or other personal qualities—but once voting begins, the electoral landscape can shift rapidly. The process is iterative: ballots are cast twice each morning and twice each afternoon, with the requirement of a two-thirds majority for a valid election. Early rounds may serve as signaling devices, allowing electors to test the waters, coalesce around emerging names, or block others (Grandori,, 2013). As such, researchers lack the empirical data needed to retrospectively analyze voting behavior or train predictive models based on past outcomes.
@@
To interpret the dynamics of the conclave that lead to the election of a pope, popular wisdom has produced a well-known saying: “He who enters the conclave as pope leaves as a cardinal.” This adage encapsulates the idea that individual attributes of a papabile represent only one component of a far more complex decision-making process, which is deeply shaped by relational dynamics. The conclave exemplifies a closed decision-making network, where each participant is both an actor with individual preferences and a node embedded in overlapping layers of influence—co-membership in dicasteries, shared consecration ties, doctrinal alignment, and regional blocs. This social structure extends beyond the cardinals admitted to the conclave itself reaching a broader set of key actors, who are part of the general congregations, consisting in preparatory meetings held in the days before the actual conclave. These meetings offer all cardinals—not only the electors—an opportunity to deliberate and prepare before voting. Unlike the conclave, which is highly codified in both procedures and rituals, the general congregations are less formalized, creating a space where existing social structures shape interaction patterns, mediation efforts, channels of influence, and the formation of preliminary agreements. As such, the boundaries of the network of relationships among cardinals that we investigate goes beyond the cardinal electors.
Understanding the conclave through the lens of social networks helps illuminate not only the institutional structure of the Church, but also the relational mechanics that drive one of the most consequential leadership selections in the modern world. The social network structure of the College of Cardinals—shaped by formal roles, shared service in Vatican bodies, and sacramental ties—provides the scaffolding for the emergence of papal frontrunners. In this context, we suggest that at least three well-established network mechanisms help explain how one candidate may progressively accumulate support and reach the two-thirds majority required for election. The first mechanism is status-based influence, which builds on the principle that connections to highly central individuals confer both visibility and endorsement potential (Podolny,, 2005). These ties generate advocacy effects: prominent cardinals can signal viability, lend symbolic weight, and help consolidate early votes behind a peer who appears not only competent and trustworthy (Soda et al.,, 2024). In a closed decision-making environment like the conclave, such cues are amplified by the limited external information available and the high interdependence among electors.
Second, individuals that stand between others in their connections (Freeman,, 1977) may facilitate or hinder the communication and information exchange, does being in a position to control the access of others to information and other resources (Freeman et al.,, 1991). By virtue of these positions, cardinals can more effectively navigate the social divides without being fully captured by any one group (Krackhardt,, 1999). In our research context, advantages do not come from re-combining information to generate novelty, but from the ability to maneuver across relational boundaries and mediate between multiple sides. Consequently, cardinals occupying these network positions can enhance their appeal as balancing candidates and gain consensus, especially in balkanized political environments (Burt,, 2010).
Third, the capacity for coalition formation often emerges from the combination of belonging to large and cohesive subgroups while at the same time reaching across the broader structure of the College. Cardinals holding bonding social capital, that is, those being part of a close-knit network, can mobilize strong and broad political support. These tightly coupled groups act as base camps for electoral coordination, facilitating early momentum and enabling wide reach as the voting unfolds. Unlike isolated factions, these groups provide both solidarity and access, increasing the likelihood that a candidate from within them can expand their support across structural boundaries.
@@
In theory, these mechanisms are not mutually exclusive. In many conclaves, the eventual winner may benefit from a combination of status, brokerage, and coalition-based advantages. A cardinal who is simultaneously connected to prominent others, positioned between distinct peers, and able to activate ties with other communities may accumulate support organically over successive ballots. In this way, the conclave’s apparent unpredictability masks a deeper process of relational convergence, driven by the structure of the electoral network itself.

### v2

# In the Network of the Conclave: Social Connections and the Making of a Pope

@@ Abstract
This study brings a network perspective to papal elections by mapping the relational architecture of the College of Cardinals. Using publicly available data sources, such as official Vatican directories and episcopal consecration records, we assemble a multiplex network that captures cardinals’ co-membership in various collegial bodies of the Vatican and their consecration ties. We then calculate structural metrics to capture three key mechanisms that we suggest have a crucial role in the dynamics of the conclave: status, mediation power, and coalition building. Our descriptive study—publicly released prior to the May 8, 2025 election of Pope Leo XIV—shows that Cardinal Robert F. Prevost, largely ignored by pundits, bookmakers, and AI models, held a uniquely advantageous position in the Vatican network, by virtue of being central in multiple respects. Thus, although being considered an underdog by many, the network perspective suggests that Cardinal Prevost was de facto one of the strongest “papabile.”
Keywords: social network analysis, centrality measures, papal conclave
@@ 1 Introduction
A: “I would not say so, in the sense that the Holy Spirit picks out the Pope… I would say that the Spirit does not exactly take control of the affair, but rather, like a good educator, as it were, leaves us much space, much freedom, without entirely abandoning us. Thus the Spirit’s role should be understood in a much more elastic sense—not that he dictates the candidate for whom one must vote.”
(Interview with Cardinal Ratzinger, later Pope Benedictus XVI)
@@
Few decision-making processes in the modern world remain as cloaked in rituals, secrecy, and symbolism as the papal conclave (Allen,, 2002). When the College of Cardinals gathers in the Sistine Chapel to elect a new pope, the world watches for the white smoke that signals the outcome of a process often described as guided by divine inspiration. Yet, as the quote above suggests, the process is not so much steered by divine forces as it is shaped by human judgment and negotiation. The fact that the Holy Spirit does not indicate which candidate one should vote for—as Cardinal Ratzinger stated—reveals that even in theological debate, it is recognized that the conclave’s decision-making process, like in many other human organizations, is not immune to mechanisms such as mediation power, coalition building, and status dynamics. Although the cardinals in the conclave are secluded from the rest of the world, they are not isolated from the force field of social influences among them built through interpersonal connections (Padgett and Ansell,, 1993; Krackhardt,, 1990). Indeed, we argue that the conclave is a complex, highly relational decision-making process, where influence flows through network ties that form the social scaffolding through which consensus can be built (Urena et al.,, 2019; Stevenson and Greenberg,, 2000). These ties create a complex network that connects the cardinals, allowing them to get to know each other, work together, exchange ideas, and transfer information (Granovetter,, 1973; Reagans and McEvily,, 2003). We ask, therefore: What role does the network—and the structural position each cardinal occupies within it—play in shaping the likelihood that someone can emerge as frontrunner candidate during the conclave?
@@
The question is as consequential as it is difficult to answer. Obtaining reliable data to understand why some cardinals emerge as key figures and how the college converges toward a candidate—and, more generally, what happens during the conclave—is, in fact, challenging, if not impossible. Those who participate in the conclave are bound to absolute silence, even years after the fact, under penalty of excommunication. [^1] The Apostolic Constitution Universi Dominici Gregis (John Paul II, 22 February 1996), which sets forth the norms for the conduct of the conclave, establishes the penalty of latae sententiae excommunication: “for those involved in the conclave who are bound to secrecy, in the event of violating that secrecy (n. 58)”.
@@
We apply social network theories and methods to enhance the understanding of the papal conclave by examining how interpersonal connections among cardinals may shape decision-making (Borgatti et al.,, 2009). This relational perspective allows us to move beyond anecdotal interpretations, offering a systematic lens for studying how patterns of connections may influence voting behavior. While certain ties—such as trust or support—may emerge or evolve during the conclave, the backbone of the network structure, built through long-standing institutional, personal, and professional relationships, originates prior to it. Analyzing the preexisting networks, therefore, provides a vantage point for anticipating the dynamics of consensus formation within the conclave. Specifically, while acknowledging that formal ecclesiastical hierarchies define roles and responsibilities, we suggest that influence and consensus dynamics within the conclave are embedded in relations, which can be made visible and analytically tractable through a network perspective. In doing so, we can model the conclave as a multiplex social system, and thus illuminate key network mechanisms underlying papal succession.
@@
Our approach is rooted entirely in network theory and mathematics, taking advantage of a decades of research on network status, information control and mediation power, and coalition building (Podolny,, 2001; Burt,, 1992; Mizruchi,, 1992). We did not leverage natural language processing, deep learning, or other artificial intelligence tools. Rather, this is a case of applying a classic social network analysis approach to a novel setting—leveraging co-membership ties and apostolic lineages—to shed light on one of the most secretive decision-making processes in the world. Our findings suggest that understanding who is papabile is not only a matter of ideology, but also of relational embeddedness, that is, being known, trusted, and connected in the right ways to the right others (Useem,, 1986; Heinz,, 1993; Krackhardt and Hanson,, 1993).
@@ 2.1 Collegial Structures
Thus, while the hierarchical structure of the Catholic Church reflects a vertical chain of command, with the pope at the top, such a depiction risks obscuring the complex, consultative, and collaboration-based processes that characterize its governance and organization. Central to this collaborative approach is the role of various collegial bodies, such as dicasteries, congregations, commissions, and councils, which support and advise the pope in exercising his office. These include institutions like the Dicastery for the Doctrine of the Faith, the Council for the Economy, and the Synod of Bishops, all of which function with a spirit of shared responsibility and collective discernment. [^4] Since the Second Vatican Council (1962–1965), there has been a significant theological and organizational emphasis on episcopal collegiality—the principle that, while the Pope holds primacy, bishops collectively share responsibility for governing the universal Church. Thus, while final authority rests with the pope, the actual process involves substantial interaction, negotiation, and consensus-building among several actors. Like in many large and complex organizations, also in the Catholic Church the extensive use of group-based decision-making and coordination mechanisms creates an articulated web of interactions among individuals by virtue of shared presence within these groups.
@@
Similar to directors sitting in corporate boards (Mizruchi,, 1996), or actors collaborating in movies (Watts and Strogatz,, 1998), the shared participation of cardinals in the formal collegial bodies of the Roman Curia constitute ties in the co-membership networks (Breiger,, 1974). Thus, to capture and analyze how structural influence unfolds among cardinals, we used a well-established technique from social network analysis by projecting a bipartite network—linking cardinals to collegial organizational bodies—into a cardinal-by-cardinal co-membership network (Pfeffer,, 2007). While our context differs, we draw methodological inspiration from Porter et al., (2005), who applied a similar approach to study co-membership patterns among legislators in the U.S. House of Representatives. In our case, the two-mode network connects cardinals to the collegial bodies to which they are assigned. Through collaboration within these bodies, cardinals cultivate mutual familiarity, exchange ideas, information, expertise, and provide reciprocal support. To summarize, the network-based lens highlights that decision-making in the Catholic Church—especially at the Vatican level—is shaped through complex webs of interlocking memberships and collaborative structures. This network originating from formal assignments constitutes a fundamental component of the salient relational infrastructure relevant to the conclave.
@@
To achieve a more fine-grained understanding of the network that connects those entrusted with electing the pope, formal ties should be complemented by more informal connections—such as friendship, advice, or trust (Krackhardt et al.,, 1992). Although our study could not reliably capture informal relationships of this kind for all cardinals, we reconstructed another type of tie that enriches the co-membership structure and helps illuminate the multiple social spaces that characterize the conclave’s decision-making process. [^5] For a small subset of cardinals, we were able to supplement our data with journalistic reports suggesting the presence of informal ties (e.g., personal friendships or long-standing collaborations). We ran our analyses both with and without these additional ties and found no meaningful differences in the results. However, because these data are not systematically available across the full sample, we do not include them in the reported network. Specifically, we focus on lines of episcopal consecration, which we describe in detail in the following section.
@@ 2.2 Consecration Lines
In the context of the conclave, these ties take on a concrete political and relational significance. A cardinal who has consecrated another member of the college—whether as principal or co-consecrator—has played a formative role in that person’s ecclesiastical trajectory. This relationship may imply a sense of loyalty, spiritual indebtedness, or alignment of vision (Yuengert,, 2001). As such, consecration lines create a relational infrastructure of influence and obligation within the electoral body, shaping how support, trust, and alignment might unfold during the papal election.
@@ 3 Network Mechanisms and the Conclave
Despite its highly codified procedures, the conclave is far from a static or predictable event. One key reason is that the actual voting preferences and dynamics remain entirely confidential, even after the end of the conclave. [^6] Allegedly, the only conclave on which there is information on the voting patterns over different rounds is the 2013 conclave that ended with the election of Cardinal Bergoglio as Pope Francis. This information comes from a book wrote by O’Connell, (2019). Ballots are burned, and no official record of individual votes is ever released. This secrecy makes each conclave a unique and analytically challenging decision-making episode. It is reasonable to expect that electors may enter the conclave with a short list of preferred candidates—often shaped by theological affinity, regional interests, curial loyalties, pastoral vision or other personal qualities—but once voting begins, the electoral landscape can shift rapidly. The process is iterative: ballots are cast twice each morning and twice each afternoon, with the requirement of a two-thirds majority for a valid election. Early rounds may serve as signaling devices, allowing electors to test the waters, coalesce around emerging names, or block others (Grandori,, 2013). As such, researchers lack the empirical data needed to retrospectively analyze voting behavior or train predictive models based on past outcomes.
@@
To interpret the dynamics of the conclave that lead to the election of a pope, popular wisdom has produced a well-known saying: “He who enters the conclave as pope leaves as a cardinal.” This adage encapsulates the historical inaccuracy of predictions, which are typically based on the individual attributes of cardinals. While such characteristics may serve as cues for voters (Berger et al.,, 1972), we theorize that the process by which cardinals become credible frontrunners is fundamentally shaped by relational dynamics. The conclave exemplifies a closed decision-making network, where each participant is both an actor with individual preferences and a node embedded in overlapping layers of influence—co-membership in dicasteries, shared consecration ties, doctrinal alignment, and regional blocs. This social structure extends beyond the cardinals admitted to the conclave itself reaching a broader set of key actors, who are part of the general congregations, consisting in preparatory meetings held in the days before the actual conclave. These meetings offer all cardinals—not only the electors—an opportunity to deliberate and prepare before voting. Unlike the conclave, which is highly codified in both procedures and rituals, the general congregations are less formalized, creating a space where existing social structures shape interaction patterns, mediation efforts, channels of influence, and the formation of preliminary agreements. As such, the boundaries of the network of relationships among cardinals that we investigate goes beyond the cardinal electors.
Understanding the conclave through the lens of social networks helps illuminate not only the institutional structure of the Church, but also the relational mechanics that drive one of the most consequential leadership selections in the modern world. The social network structure of the College of Cardinals provides the scaffolding for the emergence of papal frontrunners.
Building on this idea, we theorize that the contours of formal organizational boundaries, which in our context include relationships for coordination and resource exchange within collegial bodies, circumscribe the opportunity set for interactions to occur and develop. This concept aligns with a fundamental intuition in social network research regarding the importance of “opportunities” for social interaction, as popularized by Blau and Schwartz, (1997: 29): “…rates of social association depend on opportunities for social contact”. Within the context of organizations, as Brass and colleagues suggested (2004), people’s and groups’ formal positions constrain “…their opportunity to interact with some others and [facilitate] interaction with still others” (2004: 796). For instance, employees are more likely to interact if they share an organizational focus such as the same business unit, job function, or workspace (Kleinbaum et al.,, 2013). Thus, the design of organizational structures and processes creates the social infrastructure through which people interact. Beyond the concrete opportunities for interactions offered by the formally designed patterns, there’s also a cognitive effect generated when actors coordinate among each other within the same organizational unit. Specifically, research showed that organizational units identify discursive entities or “zones of meaning that are linguistically circumscribed” Tasselli et al., (2020: 1294). Therefore, membership in groups within an organization, entails both adopting and influencing the shared interpretations that individuals employ to comprehend the organization’s social landscape. Thus, formal organizational structures influence organizational actors in developing shared cognitive schemas—particularly among those who belong to the same units—that pertain to various factors: the social structure of the organization, the subgroups within it, the status hierarchy that defines relationships among those subgroups, and the roles individuals play within them (Balogun and Johnson,, 2005; Podolny,, 2001).
In this context, we suggest that at least three network mechanisms help explain how one candidate may progressively accumulate support and reach the two-thirds majority required for election. The first mechanism is status-based influence, which builds on the principle that connections to highly central individuals confer both visibility and endorsement potential (Podolny,, 2005). These ties generate advocacy effects: prominent cardinals can signal viability, lend symbolic weight, and help consolidate early votes behind a peer who appears not only competent and trustworthy (Soda et al.,, 2024). Ties to high-status actors do not merely boost a cardinal’s prestige through association, they actively enhance perceptions of his competence and reliability, thereby increasing his perceived trustworthiness and solidifying the candidacy. Additionally, since the relational infrastructure underpinning the network of cardinals is the outcome of specific organizational choices, prominence within this network also serves as an institutional signal, a normative cue that may further shape voting behavior. Cardinals are aware of each other’s roles, affiliations, and histories and the network in which they are embedded formed by co-membership ties in the collegial bodies and the consecration lines is largely observable from within. In highly symbolic settings such as the conclave, decisions are shaped by normative roles and institutional validation. Under this logic, the candidates’ credibility hinges on whether their affiliations resonate with the institutional logic, that is, the implicit repertoire of what is considered “appropriate” within the conclave’s social and historical context. Thus, what matters is not merely competence in isolation, but whether a candidate’s identity and bearing are perceived as fitting the conclave’s normative fabric—an identity that adheres to the collective understanding of what a pope should embody.
Second, individuals who stand between others in their connections (Freeman,, 1977) may facilitate or hinder the communication and information exchange by occupying positions that allow them to control others’ access to information and resources (Freeman et al.,, 1991). By virtue of these positions, cardinals can more effectively navigate the social divides without being fully captured by any one group (Gargiulo and Benassi,, 2000), allowing them to influence perceptions and manage tensions (Burt,, 1992). Cardinals occupying these positions can bridge ideological, geographical, or institutional divides without appearing fully aligned with any single faction. This positional flexibility enables such candidates to be perceived as neutral integrators, particularly appealing in politically fragmented or polarized settings (Gould and Fernandez,, 1989). Unlike contexts in which mediators and brokers derive value from recombining diverse information to foster innovation (Burt,, 2004), the conclave may reward these positions since they help in connecting boundaries by maintaining cohesion across camps. Their ability to interpret, reframe, or transmit messages across possible divides makes them instrumental in fostering consensus. This bridging capacity enhances their symbolic legitimacy as unifying figures, positioning them as "compromise candidates" or stabilizing choices in times of institutional ambiguity (Burt,, 2010).
Third, the capacity for coalition formation often hinges on the interplay two network properties (Reagans and McEvily,, 2003): deep integration within cohesive subgroups and simultaneous reach across the broader structure of the College of Cardinals. This mechanism reflects the interplay between bonding and bridging social capital (Adler and Kwon,, 2002). Candidates embedded in dense, tightly-knit networks benefit from bonding capital, which provides trust, mutual obligation, and coordinated action. These close-knit groups function as electoral base camps, where early support can be consolidated, communication is rapid and reliable, and strategic coordination is facilitated (Coleman,, 1988). Such structure plays a crucial role in the early rounds of voting, where uncertainty is high and the field of contenders still fluid. The internal solidarity of these subgroups enables them to act as cohesive voting blocs, amplifying their influence by signaling credibility and organizational capacity to the rest of the electorate. In this way, candidates from within these groups gain an initial momentum advantage, which can be essential in shaping perceptions of viability and setting the tone for broader coalition-building. At the same time, the ability to extend relational ties beyond the subgroup—that is, to engage across factional or regional divides—is what transforms a local base into a platform for wide-reaching support. Whereas isolated factions may foster loyalty, their insularity limits the candidate’s capacity to scale influence. By contrast, candidates who can anchor themselves in cohesive communities while also maintaining outward-facing ties are uniquely positioned to orchestrate cross-cutting alliances (Burt,, 2005; Obstfeld,, 2005). Their social position enables them to translate local solidarity into broader legitimacy, enhancing their appeal as coalition candidates capable of unifying diverse constituencies within the College. In sum, this dual network configuration strengthens not only a candidate’s organizational support base but also their symbolic capacity to unify, making them especially effective in navigating the political complexity of conclave dynamics.
@@
In theory, the candidate who succeeds may benefit from a combination of status, brokerage, and coalition-based advantages. Specifically, a cardinal who is simultaneously connected to prominent others, positioned between distinct peers, and able to activate ties with other communities may accumulate support organically over successive ballots. In this way, the conclave’s apparent unpredictability masks a deeper process of relational convergence, driven by the structure of the electoral network itself.